OpenStudy (anonymous):

Quick question (checking) in math: \[\text{If we have }\frac{x}{x-3}-\frac{1}{x+1}\]\[\text{ then how will we simplify after simplifying to}\frac{x(x+1)-(x-3)}{x^{2}-2x-3}\]

OpenStudy (anonymous):

Basically I'm not sure whether to go\[\frac{x^{2}+x-x+3}{x^{2}-2x-3}\]or to this:\[\frac{x^{2}+x-x-3}{x^{2}-2x-3}\]

OpenStudy (anonymous):

Since the parentheses doesn't matter... does it?

OpenStudy (bibby):

it does. you're subtracting the quantity. you can th ink of it as distributing the negative. \(x^2+x-(x-3)=x^2+x-x--3=x^2+3\)

OpenStudy (anonymous):

ok thanks